Assuming that the perpetual inventory method is used and costs are computed at the time of each withdrawal, what is the gross profit if the inventory is valued at FIFO?
Answer #1
Total sales = (199*24)+(497*26)+(295*28)= $25958
Cost of goods sold = (298*10)+(693*12)= $11296
Gross Profit = Sales-Cost of goods sold = 25958-11296= $14662